Circular No. 7610

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2000 (1998 WW_{31}) 1

C. Veillet, Canada-France-Hawaii Telescope (CFHT), reports: “Recovery images of the transneptunian object (TNO) 1998 WW_{31} (see MPEC 2001-G29) taken by C. Veillet, A. Doressoundiram, and J. Shapiro with the 3.6-m CFHT show that two objects were within < 1′′.3 and moving together over the two nights of observations (2000 Dec. 22 and 23 UT) without any detectable relative motion. CFHT public archive observations of the same field taken nearly a year previously by J. J. Kavelaars and A. Morbidelli show 1998 WW_{31} as double or elongated on four images, with the two components at a different distance and position angle than on the discovery images. A very preliminary reduction shows the brighter component to be 0.4 mag brighter than the secondary in R on 2000 Dec. 22, with the secondary being 1′′.2 from the primary in p.a. 45^{o} (with seeing 0′′.7–1′′.1). The archival images taken on 2000 Jan. 7.3 show the secondary to be 0′′.8 from the primary in p.a. 25^{o}. Images taken on 2000 Jan. 6 show the same elongation, but the pair is embedded in a bright star halo, making any measurement difficult. The maximum distance between the two components is thus at least 40 000 km. This indicates that 1998 WW_{31} is the second TNO to have a satellite (after Pluto). The analysis of other images from the CFHT, from Kitt Peak (1998 Nov. 18, 1999 Jan. 14, 2000 Nov. 23), and from the Nordic Optical Telescope (1998 Dec. 18) may allow a full orbit determination, leading to physical parameters of the pair. Images and details on the data are

available at http://cfht.hawaii.edu/~veillet/WW31.html.”

V382 VELORUM M. Bos, Mount Molehill Observatory, Auckland; A. Retter, Keele Uni-

versity; and J. McCormick and F. Velthuis, Farm Cove Observatory, Auckland, report: “Unfiltered CCD photometry of N Vel 1999 on twelve nights during Jan.-Mar. reveals the presence of a periodicity of 0.14615(1) day in the light curve of this classical nova together with its first harmonic. The semiamplitude of the variation is 0.02 mag.”

Visual magnitude estimates by A. Pearce, Nedlands, W. Australia: 2000 May 2.59 UT, 10.3; Sept. 23.866, 11.6; Dec. 2.767, 11.8; 2001 Feb. 26.788, 12.3; Mar. 5.788, 12.4.

2001 April 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green